OpenStudy (anonymous):

tan a = 5/12 tan b = root3/3 tan (a-b)?

OpenStudy (anonymous):

\[\frac{\tan(\alpha)-\tan(\beta)}{1+\tan(\alpha)\tan(\beta)}\] substitute the numbers directly in to the formula

OpenStudy (anonymous):

I ended up with (5/12)-(4root3/12)/1+5root3/36?